OpenStudy (russianboss1):

Fan and Medal!!A parabola had vertex(3,2) and contains the point(-1,14). What is the equation of the function.?

OpenStudy (russianboss1):

@gottennis121

OpenStudy (gottennis121):

Well, we know what one of the points are, -1, 14. We also know that parabola (the bottom of the upside down "u" shaped figure) as 3,2. Is this multiple choice?

OpenStudy (gottennis121):

@lsherron

OpenStudy (gottennis121):

a)y=3/4(x-3)^2+2 b)y=2/3(x-3)^2+2 c)y=3(x-3)^2+2 d)y=(x-3)^2+2

OpenStudy (anonymous):

\[f(x)=a(x-3)^2+2\] put \(x=-1\) set it equal to \(14\) and solve for \(a\)

OpenStudy (anonymous):

in other words, solve \[14=a(-1-3)^2+2\] for \(a\)